### v1

# Symmetrical Visual Contrastive Optimization: Aligning Vision-Language Models with Minimal Contrastive Images

Abstract
Recent studies have shown that Large Vision-Language Models (VLMs) tend to neglect image content and over-rely on language-model priors, resulting in errors in visually grounded tasks and hallucinations. We hypothesize that this issue arises because existing VLMs are not explicitly trained to generate texts that are accurately grounded in fine-grained image details. To enhance visual feedback during VLM training, we propose S-VCO (Symmetrical Visual Contrastive Optimization), a novel finetuning objective that steers the model toward capturing important visual details and aligning them with corresponding text tokens. To further facilitate this detailed alignment, we introduce MVC, a paired image-text dataset built by automatically filtering and augmenting visual counterfactual data to challenge the model with hard contrastive cases involving Minimal Visual Contrasts. Experiments show that our method consistently improves VLM performance across diverse benchmarks covering various abilities and domains, achieving up to a 22% reduction in hallucinations, and significant gains in vision-centric and general tasks. Notably, these improvements become increasingly pronounced in benchmarks with higher visual dependency. In short, S-VCO offers a significant enhancement of VLM’s visually-dependent task performance while retaining or even improving the model’s general abilities.
Symmetrical Visual Contrastive Optimization: Aligning Vision-Language Models with Minimal Contrastive Images
Shengguang Wu, Fan-Yun Sun, Kaiyue Wen, Nick Haber Stanford University https://s-vco.github.io/
1 Introduction
Large Vision-Language Models (VLMs) tend to over-rely on their language models, leading to neglect of visual content. This problem manifests as visual hallucinations across tasks like perception (Tong et al., 2024b), reasoning (Chen et al., 2024), and in-context learning (Jia et al., 2024). Studies like Tong et al. (2024a) show that VLMs exhibit only limited performance gains when vision inputs are enabled compared to having no vision inputs. This behavior is reflected in the metrics across many popular vision-language benchmarks (Lu et al., 2022; Yue et al., 2024; Lu et al., 2023; Kembhavi et al., 2016; xAI, 2024; Singh et al., 2019). Our own perplexity-based evaluations reveal similar patterns of visual neglect. As illustrated in Figure 2, we measured a base VLM’s perplexity (PPL) when generating a caption matched to one image (e.g., “imagematch” with a “dog”) while contradicting the other image (e.g., “imagemismatch” with a “cat”). We also tested the model’s PPL without any vision input (“no image”). Intuitively, the model should exhibit the lowest PPL when given the matching image, as the aligned visual information would make the caption easier to generate. However, results reveal the opposite pattern: the model’s PPL lowest when no image input is provided at all, and highest when presented with the correct image (the “dog”) – higher than the mismatched image (the “cat”) as condition. Across random samples from CounterCurate (Zhang et al., 2024) — a large collection of such paired counterfactual images — the average perplexity distribution follows this counterintuitive pattern. This highlights the model’s tendency to disregard visual information even when they are critical for generating accurate texts.
Recent works (Wang et al., 2024; Xie et al., 2024; Jiang et al., 2024a; Luo et al., 2024; Fu et al., 2025) attempt to address similar issues of visual neglect by adapting Direct Preference Optimization (DPO) (Rafailov et al., 2024) to compare image inputs. However, these methods treat visual supervision as a “preferential” tuning paradigm, where an original image is preferred, and a cropped or noisy version of that image is dispreferred. As shown in Figure 3, these noisy images lack meaningful connections to their paired texts, making it easy for the model to learn shortcuts by rejecting the corrupted image versions without fully understanding visual details and aligning them to the corresponding texts.
Instead, we posit that the “preference” paradigm should be improved with a purely contrastive framework, as the “dispreferred” image is merely another image misaligned with a given text. Building upon this insight, we propose Symmetrical Visual Contrastive Optimization (S-VCO), a finetuning objective that enforces precise correspondence between visual details and textual tokens. S-VCO rewards the model for attending to matching images and strongly rejecting contradictory images with incorrect details. To further avoid shortcut learning, S-VCO incorporates symmetry by flipping the objective for contradictory responses, allowing the “negative” image to serve as the “preferred” visual condition when paired with its corresponding text.
Additionally, as cropping Wang et al. (2024) or adding diffusion noise to an original image Jiang et al. (2024a) fails to provide meaningful comparisons in visual details, we construct MVC, a dataset of paired images with Minimal Visual Contrasts, each matched to contrastive textual responses given a shared query. Building on recent visual counterfactual data sources (Zhang et al., 2024; Liu et al., 2024c), we implement a vision-centric filter to select visually challenging pairs and an LLM augmentation scheme to diversify texts, forming an instruction-response-styled dataset suitable for VLM finetuning.
Experiments demonstrate the effectiveness and versatility of our approach, as S-VCO consistently enhances VLM performance across diverse benchmarks spanning various abilities and domains. As shown in Figure 1, S-VCO achieves significantly greater improvements over the base-VLM, particularly in reducing visual hallucinations, while excelling in vision-centric tasks and offering considerable gains on general benchmarks. Compared to existing VLM preference tuning methods (DPO, Rafailov et al., 2024; mDPO, Wang et al., 2024), S-VCO combined with MVC delivers more substantial and comprehensive performance boosts.
In summary, our main contributions are:
• S-VCO, a novel VLM finetuning objective that enforces strict and balanced visual contrastive supervision through the symmetrical alignment of image-text pairs;
• MVC, a dataset of minimal contrastive image pairs accompanied with corresponding textual responses for a shared query. MVC is constructed automatically through vision-centric filtering and augmentation based on visual counterfactual data;
• The combination of S-VCO and MVC significantly boosts VLM performance across diverse benchmarks, particularly in visually dependent tasks, without compromising general capabilities.
2 Preliminaries
Our visual contrastive objective is inspired by Direct Preference Optimization (DPO) (Rafailov et al., 2024) that contrasts pairs of textual responses. When applied to VLMs, DPO incorporates the image as an additional prefix condition (Zhou et al., 2024; Li et al., 2023a; Yu et al., 2024; Sarkar et al., 2024). Let be the policy VLM to be optimized, and the fixed reference model (typically the unfinetuend VLM in DPO’s framework) used to measure and optimize how the finetuned policy improves. Given a query (an instruction prompt), an image , and a pair of responses – one preferred (winning) and one dispreferred (losing) – the DPO objective for VLMs can be formulated as:
As illustrated in Figure 3, the DPO formulation focuses on supervising differences between language responses, which however, often pertain to wording choices and stylistic variations rather than reflecting meaningful visual distinctions. Consequently, the VLM is being trained like a language model to weigh over text formulations, rather than to fully utilize the image as a grounding input.
Recent approaches, such as Wang et al. (2024); Jiang et al. (2024a), adapt the DPO framework to contrast image conditions. The original image becomes the preferred image (winning), while a negative image (losing) is derived through random cropping (Wang et al., 2024) or adding diffusion noise (Jiang et al., 2024a) (see Figure 3). The updated Visual Conditional objective is:
The above visual conditional preference formulation shifts the target of preference to the images, where the original image is always preferred over the corrupted version . However, this approach shares similar limitations with DPO, as it prioritizes distinguishing image differences without necessarily grounding those differences in the associated texts. Since is typically a noisy variant of with no definitive relationship to , the model could easily rely on superficial visual features to discern the paired images. This encourages shortcut learning, where the model rejects “unrealistic” images like without examining the visual details related to the text tokens.
Our S-VCO addresses these limitations by introducing a stricter visual-conditioned objective (Section 3.1) and a symmetrical construct that aligns both and with their respective and (Section 3.2). Figure 3 illustrates S-VCO in comparison to existing preference tuning paradigms, showcasing its emphasis on grounded visual-textual alignment.
3 S-VCO: Symmetrical Visual Contrastive Optimization
3.1 Visual Contrastive Supervision
S-VCO enforces a strict visual focus by optimizing the model for two key behaviors:
1) Attending to matching images. The model is rewarded for prioritizing relevant visual details in the matching image as a condition when predicting the corresponding response . This directly addresses the tendency of VLMs to overlook visual content (Section 1) and is achieved through the term:
2) Rejecting contradictory images. When presented with a contrastive image containing visual details that directly contradict the response , the model must strongly reduce the likelihood of predicting under this incorrect image condition. Intuitively, the model should assign minimal probability to a response that directly contrasts with the visual input. This behavior is modeled as:
By combining these two components, our strict visual contrastive objective is defined as:
3.2 Symmetrical Alignment
Unlike prior “preference”-based approaches, S-VCO treats and as mere images with contrastive details, where either can serve as the correct (i.e., “preferred”) condition depending on the paired textual response ( or ). While we inherit the notation of and , S-VCO does not assign an inherent “winning” or “losing” property to the images. Instead, an image is considered “winning” only when paired with its corresponding text. For instance, , typically labeled as “losing” in preference tuning methods, becomes a “winning” (preferred) condition when the target response is that matches its visual details.
A one-sided formulation that consistently favors over risks encouraging shortcut learning, where the model rejects based on superficial, text-unrelated features (e.g., visual style differences). This issue could arise because most of the contrasting images are synthesized via inpainting or image editing (see Section 4). To address this, S-VCO introduces symmetry by flipping the objective in Equation 5, treating as the preferred condition when paired with its corresponding response . This effectively switches the roles of “winning” and “losing” for the image pair:
As defined above, this encourages the model to attend to and reject in the flipped case where the target response is .
The complete S-VCO objective incorporates symmetry by summing over both roles of and given their corresponding target responses:
This symmetrical alignment ensures balanced optimization, allowing both and to contribute equally to the model’s learning. It promotes true alignment between images and texts without relying on shortcuts, such as rejecting either the image (Equation 2) or the text (Equation 1). In essence, S-VCO optimizes for the alignment of image-text pairs rather than simply one modality.
4 Minimal Visual Contrasts Dataset
4.1 Visual Counterfactual Data
To complement S-VCO, we introduce MVC: a dataset of paired visual contrastive samples designed to enhance the model’s ability to discern visual details. Built on existing sources (e.g., Zhang et al., 2024; Liu et al., 2024c; Gaur et al., 2024), MVC contains image pairs with minimal but meaningful variations, accompanied by corresponding contrastive texts. The curated data includes four key contrast types, as shown in part a). of Figure 4: Object Replacement (changing a specific object); Attribute Replacement (modifying an object’s features like color, shape, or size); Count Modification (altering the number of objects); Position Flipping (reversing relative positions of objects). Except for the last type, these counterfactual images are generated through controlled image synthesis, including inpainting, editing, and generation (Li et al., 2023b; Zhuang et al., 2025; Betker et al., ). In this work, we use CounterCurate (Zhang et al., 2024) and FineCops-Ref (Liu et al., 2024c) as our visual counterfactual data sources. Refer to Appendix A for more dataset statistics.
4.2 Filtering and Language Augmentation
Filter. Existing visual counterfactual datasets offer a large quantity of detailed contrasts, but their quality is inconsistent due to the synthetic nature of the data, which could lead to pairs where the generated image fails to truly contradict the original. To address this, we implement a vision-centric filter inspired by the CLIP-Blind concept (Tong et al., 2024b) to select image pairs based on the following two criteria, as shown in part b). of Figure 4: 1. Different in detailed visual features: Image pairs must exhibit meaningful contrasts, especially in detailed visual features. To achieve this, we employ DINOv2 (Oquab et al., 2023), a vision-only model with more robust visual feature representations (Singh et al., 2023). Pairs with high similarity in DINOv2’s purely visual representation space are discarded, as they may not contain the desired degree of contrasts. 2. Semantically close & Hard for VLMs: Image pairs must also be semantically similar overall and difficult for current VLMs to distinguish. Therefore, we embed images using the same CLIP vision encoder used by the VLM (Radford et al., 2021; Zhai et al., 2023) and retain pairs with relatively high similarity in the CLIP space that focuses on images’ overall visual semantics. In this way, we exclude pairs with overly distinct content and less significant contrasts in visual details. In practice, we use DINOv2-Large (Oquab et al., 2023) with a similarity threshold of , and SigLIP-400M (Zhai et al., 2023) as the CLIP encoder with a similarity threshold of . Together, these thresholds ensure that MVC focuses on meaningful visual contrasts while maintaining semantic relevance and difficulty for the VLM.
Language Augmentation. Although visual counterfactual data sources provide image contrasts, their original textual descriptions are typically short captions, which are unideal for VLM finetuning. To address this, we augment the data using a two-step process with a strong LLM (gpt-4o), as shown in part c). of Figure 4: 1. Generating queries: For each pair of captions, we prompt the LLM to generate a conversational question as if the captions were natural responses to that question, while ensuring the contrasts remain explicit. 2. Rewriting and diversifying: We prompt the LLM to rephrase the original captions lacking clarity on key contrasts, thus enabling a more natural flow of language given the generated question and the contrastive details. Overall, the augmentation step results in conversational instruction-response pairs that are more suited for VLM finetuning, and more aligned with the visual details in the contrasting image pairs. Our prompt template is provided in Appendix C.
After the filtering and language augmentation, our MVC dataset comprises over pairs of minimal contrastive images matched with accurate and diverse conversational queries and responses (Appendix A). The effectiveness of these data processing steps is empirically discussed in Section 5.3.
| Benchmarks | Hallucination | Vision-Centric | General | OCR | Knowledge | TOTAL | |||||
| MMHal | CVBench | MMVP | RQA | MMVet | LVBench | TextVQA | SQA | avg_impr. | |||
| Models | score | hal_rate | acc. | acc. | acc. | score | score | acc. | acc. | over BASE | |
| LV-1.5-7B | BASE | ||||||||||
| DPO | |||||||||||
| DPO | |||||||||||
| mDPO | |||||||||||
| mDPO | |||||||||||
| S-VCO | |||||||||||
| LV-INT-7B | BASE | ||||||||||
| DPO | |||||||||||
| DPO | |||||||||||
| mDPO | |||||||||||
| mDPO | |||||||||||
| S-VCO | |||||||||||
| S-VCO | |||||||||||
| VCO | |||||||||||
| SFT | |||||||||||
5 Experiments
5.1 Setup
Baseline methods. We compare S-VCO against DPO (Rafailov et al., 2024) and mDPO (Wang et al., 2024)111In practice, mDPO combines textual DPO (Equation 1) with the visual-conditional version (Equation 2) and an absolute reward regularization. containing the visual-conditional objective in Equation 2 as baseline finetuning approaches.
Training data. We use two datasets: our proposed MVC with minimal contrastive image-text pairs and VLFeedback (VLF) (Li et al., 2023a), a classical instruction-tuning dataset for VLMs that includes preferred and dispreferred response pairs. We follow mDPO (Wang et al., 2024)’s sample of ~10,000 data points from VLF. Refer to Appendix A for more dataset implementation details.
Base VLMs. We use two pretrained models hosted on Huggingface: LLaVA-1.5-7B222llava-hf/llava-1.5-7b-hf (Liu et al., 2024a) – denoted as LV-1.5(-7B), and LLaVA-Next-Interleave-7B333llava-hf/llava-interleave-qwen-7b-hf (Li et al., 2024) – denoted as LV-INT(-7B).
Evaluation. We evaluate the models on a wide range of benchmarks spanning various ability domains. Following the categorization by Tong et al. (2024a), these benchmarks include: General: LLaVABench (Liu et al., 2024b), MMVet (Yu et al., 2023); Hallucination: MM-Hal (Sun et al., 2023); Vision-Centric: CVBench (Tong et al., 2024a), MMVP (Tong et al., 2024b), RealworldQA (xAI, 2024); OCR: TextVQA (Singh et al., 2019); Knowledge: SQA (Lu et al., 2022).
Implementation details. Refer to Appendix B for detailed training and inference configurations.
5.2 Main Results
Table 1 presents detailed evaluation results for different methods across various benchmarks grouped by ability domains. The final column shows the average improvement (%) over the base VLM across all metrics, summarizing each model’s overall performance. Key findings are highlighted below.
S-VCO achieves consistently superior performance across benchmarks in various domains. Figure 1 illustrates performance improvements over the base-VLM (LV-INT) for each benchmark category. Our S-VCO consistently outperforms baseline methods across nearly all domains, with only a slight drop in knowledge-heavy tasks like ScienceQA that relies minimally on visual input (discussed more below). The most significant gain is in visual hallucination tasks, where S-VCO enhances the base model by over %. Considerable improvements are also observed in vision-centric (%) and general domains (%). Compared to baseline methods, DPO and mDPO trained on the standard instruction-tuning dataset VLF, S-VCO with MVC delivers much greater and more consistent improvements across domains. While recent approaches like mDPO (Wang et al., 2024) improve visual hallucination metrics, its effects on other abilities are less notable. In contrast, S-VCO not only excels in visually demanding tasks but also achieves considerable gains in other domains.
S-VCO shows increasing benefits as benchmarks become more visually dependent. We quantify visual dependency of a metric as the percentage drop in a base-VLM’s performance when image inputs are removed. Using LV-INT as the base model, we rank benchmarks by visual dependency and plot the improvement trends of different methods in Figure 6. Dotted lines show fitted regressions, and shaded areas represent variances. S-VCO demonstrates increasingly pronounced improvements as visual dependency rises, aligning with its design focus on strengthening visual detail recognition (Section 3). For visually highly dependent benchmarks like LLaVABench (Liu et al., 2024b), MMVet (Yu et al., 2023), and MM-Hal (Sun et al., 2023), S-VCO delivers the most substantial gains. In contrast, ScienceQA (SQA) (Lu et al., 2022), which benefits merely ~% from image inputs, sees a minor drop in performance, as visual information plays little role in this task. Compared to other methods, SFT (Section 5.3) degrades performance on visually dependent metrics, while other preference tuning methods show positive trends. Among all, S-VCO achieves the most significant trend of gains with increasingly vision-intensive tasks.
Qualitative performance highlights S-VCO’s superior visual understanding. In Figure 5, qualitative examples from various benchmarks further demonstrate S-VCO’s superior ability to process fine-grained visual details and reason about complex scenes. S-VCO excels in recognizing subtle yet critical visual distinctions (e.g., identifying the absence of a toothbrush) and remains robust against visual hallucinations (e.g., differentiating marker-drawings, slide-phones, and fire hydrants). Moreover, S-VCO shows strong visual reasoning capabilities by interpreting complex scenarios such as drive-lane conditions. It also captures intricate details in scenes with greater depth and contextual awareness (e.g., discerning weather conditions through a window or identifying oncoming vehicles in low-light settings).
MVC enhances previous preference tuning methods. Beyond its strong synergy with S-VCO, MVC dataset also strengthens the effects of existing preference tuning methods, particularly DPO. For LV-INT, both DPO and mDPO achieve greater overall improvements when trained on MVC compared to the textual-instruction-tuning-styled dataset VLF. For LV-1.5, DPO outperforms DPO by ~% on average across benchmarks, achieving second-best results on multiple metrics behind only our S-VCO. Figure 6 further illustrates DPO and mDPO variants trained on MVC outperforming their VLF-trained counterparts, especially on visually dependent tasks. By providing visually challenging image pairs matched with texts, MVC proves to be a more effective training resource across preference tuning methods.
5.3 Ablations
Data filter and augmentation. We evaluate the impact of our filtering and augmentation step (Section 4.2) in MVC by training LV-INT-7B on unprocessed (“raw”) visual counterfactual data directly sourced from CounterCurate (Zhang et al., 2024) and FineCops-Ref (Liu et al., 2024c). We sample random datapoints exactly matching MVC’s size (Appendix A), and train the model with S-VCO using the same configurations. The resulting model, S-VCO, underperforms S-VCO trained on filtered and augmented data, with drops across most benchmarks (Table 1). This highlights the importance of our filter and augmentation step in constructing high-quality contrastive data that better supports S-VCO. Notably, despite the data preprocessing omissions, S-VCO still surpasses all other baseline preference tuning methods, demonstrating the robustness of S-VCO’s symmetrical visual contrastive objective.
Symmetrical loss construction. A key feature of S-VCO is its symmetrical loss, which treats both sides of the texts and as preferred when aligned with their respective images and , optimizing both simultaneously (Section 3.2). To assess the necessity of this symmetry, we train LV-INT-7B on MVC without the symmetrical term , reducing the objective to a one-sided preference tuning approach. The resulting model, VCO, underperforms S-VCO, as shown in Table 1, though it still surpasses all other baseline methods due to VCO’s strong visual contrastive supervision (Section 3.1). The performance drop is most evident in hallucination and vision-centric tasks, highlighting the importance of optimizing both sides of the contrastive pair. Symmetry mitigates shortcut learning by encouraging the model to focus on meaningful image-text alignments rather than superficial image features.
Comparing to SFT. To investigate whether standard supervised finetuning (SFT) on both sides of the image-text pairs could achieve similar results, we construct a new instruction-tuning dataset by including both and from MVC, effectively doubling its size (). Results for SFT are shown in the last row of Table 1. While SFT preserves performance on ScienceQA – a benchmark minimally reliant on visual inputs – it performs significantly worse across all other domains, achieving the lowest results among all methods. Unlike S-VCO, SFT lacks the contrastive supervision necessary to highlight subtle visual-textual alignments, thus leading to poor performance on tasks requiring strong visual grounding.
6 Related Work
VLM’s Visual Hallucinations. Recent studies (Dai et al., 2024; Deng et al., 2024; Tong et al., 2024b; Chen et al., 2024; Jia et al., 2024) observed the tendency of VLMs to hallucinate content that is not present in the visual input, indicating a lack of robust multimodal grounding. To address this issue, several training-free methods have been proposed. Yang et al. (2023) introduced the “Set-of-Mark” prompting technique, which overlays spatial and textual markers on images, helping models like GPT-4V better reference specific regions. Deng et al. (2024) employed CLIP-guided decoding to steer the language outputs with grounded visual cues. On the architectural side, GRILL Jin et al. (2023) incorporates object-level alignment during pretraining to prompt visual grounding. In contrast to these approaches, our work focuses on finetuning with a novel preference-tuning objective (Section 3) and a data construction pipeline (Section 4) based on visual counterfactuals (Zhang et al., 2024; Liu et al., 2024c), targeting more precise alignment between multimodal details.
VLM Finetuning. Finetuning enhances VLMs for task-specific performance and alignment with human preferences. SFT remains widely adopted to guide models toward instruction-following behaviors in multimodal contexts (Sun et al., 2024; Jiang et al., 2024b). DPO (Rafailov et al., 2024) optimizes the margin between the finetuned and unfintuned model versions using paired preference data. Extensions of DPO to VLMs incorporate the image as an additional prefix condition (Zhou et al., 2024). Recent methods such as mDPO (Wang et al., 2024), MFPO (Jiang et al., 2024a), V-DPO (Xie et al., 2024), CHiP (Fu et al., 2025) and Image-DPO (Luo et al., 2024) further adapt the preference tuning paradigm to focus on image-side preferences over a pair of “good” and “bad” image, aiming to reduce visual hallucinations. Our approach S-VCO discards the one-sided “preference” formulation by introducing a stricter visual contrastive objective within a symmetrical construct, in which way the “preference” is treated as alignment over a matching image-text pair. This enables more comprehensive and robust improvements in VLM performance across tasks.
7 Conclusion
This work introduces S-VCO, a novel VLM finetuning objective that enforces strict visual contrastive supervision within a symmetrical construct. Complementing this objective, we propose MVC, a dataset of paired images with minimal visual contrasts, each associated with corresponding contrastive texts. Experiments demonstrate that combining S-VCO with MVC consistently improves VLM performance across diverse benchmarks, with particularly significant gains on visually dependent tasks. Importantly, these improvements are achieved without compromising, and even enhancing VLMs’ general capabilities.
Limitations
Our method incorporates multiple individual loss terms and weights (Section 3), which may require manual tuning to determine the optimal settings. Adjusting these hyperparameters could potentially further enhance performance.
The S-VCO objective works most effectively when the contrastive image pairs have meaningful differences in visual details. The current MVC derived from existing visual counterfactual data sources includes a limited set of operations for building the contrasts (Section 4.1). Our method could benefit from a more diverse set of contrasting visual details that should enable the model to potentially learn a broader range of visual features.
Ethics Statement
In this work, all data and pretrained models are publicly available. They are collected and processed in adherence to the respective data, checkpoints, and API usage policy. We acknowledge that our finetuned models may generate unsafe content, and we advise all users of careful verification before deploying this work in real-world applications.
Appendix A Dataset Details
Table 2 details the dataset statistics of our visual counterfactual data sources CounterCurate (Zhang et al., 2024) and FineCops-Ref (Liu et al., 2024c), as well as our MVC after filtering and augmentation (disccused in Section 4.2). We discard the original “Order” category of FineCops-Ref for the frequent irrational cases under that category. For the non-synthetic “Left-Right” position flipping, we did not apply the filter but randomly sampled the data proportional to its original category distribution in the source datasets.
When training with MVC, we leave out random samples for validation set ( from CounterCurate and from FineCops-Ref). This leads to a total training data size of with MVC. When training with the sampled VLFeedback data (Li et al., 2023a), used as in mDPO (Wang et al., 2024) (https://huggingface.co/datasets/fwnlp/mDPO-preference-data; noted as VLF), we leave out random samples for validation set, leading to a total training data size of with VLF.
| CounterCurate | FineCops-Ref | MVC | |
| Object Replacement | |||
| Attribute Replacement | |||
| Count Modification | |||
| Position Change | |||
| Total |
Appendix B Implementation Details
Training. We set and in S-VCO’s objective (Equation 7) to , following the typical values in DPO (Equation 1) and mDPO (Equation 2). All models are finetuned for epochs with a batch size of on NVIDIA-A100x80G GPUs. When training on VLF, we retain the original learning rate of used by DPO and mDPO. The text sequence length during finetuning is set to to accommodate VLF’s text data length. When training on our MVC, we set a learning rate of for all methods except mDPO on LV-1.5-7B, where a lower rate of is used to better stabilize training. The text sequence length during finetuning is set to given the MVC’s shorter text data length. Intermediate checkpoints are saved at an interval of , , and steps for training on VLF, MVC, and MVC (used for SFT ablation, see Section 5.3), respectively.
Evaluation. We report results using the best-performing checkpoint for each model configuration (Table 1), selected based on the highest average improvement over the base-VLM. For all benchmarks, the temperature of model predictions is set to . As the evaluation judge, we use gpt-4-0613 for LLaVABench (Liu et al., 2024b) and MMVet (Yu et al., 2023), and gpt-4-turbo for MM-Hal (Sun et al., 2023).
Appendix C Prompt for MVC Language Augmentation
Below are our prompt templates for querying GPT4 (gpt-4o) to augment the queries and responses from visual counterfactual data sources (Section 4.2).
In the first step, we ask GPT4 to generate an natural and appropriate question that targets the subtle differences in the response-pair.
In the second step, we ask GPT4 to revise the instruction generated from the first step, and rephrase the original response pairs to make the whole instruction-response conversation sound more natural, while retaining the contrastive details in the responses.
In both steps, we set the temperature to for more diversified wording.
[h]
[h]